May you please help me with this question? Which is:
Q:A rectangular pool measures 12 meters by 8 meters and is bordered by a concrete path 2 meters wide.
How many square meters are in the surface area of the concrete path?
Thank you

Answer by rothauserc(4718)

A rectangular pool measures 12 meters by 8 meters and is bordered by a concrete path 2 meters wide.
How many square meters are in the surface area of the concrete path?
==============================================================================
we want to subtract the inner rectangle's area from the larger rectangle's area
surface area of concrete path = 16*12 - 12*8 = 192 - 96 = 96 square meters